# Practical Collaborative Perception: A Framework for Asynchronous and Multi-Agent 3D Object Detection

### v1

@@ Abstract
In this paper, we improve the single-vehicle 3D object detection models using LiDAR by extending their capacity to process point cloud sequences instead of individual point clouds. In this step, we extend our previous work [1] on rectification of the shadow effect in the concatenation of point clouds to boost the detection accuracy of multi-frame detection models. Our extension includes incorporating HD Map and distilling an Oracle model. Next, we further increase the performance of single-vehicle perception using multi-agent collaboration via Vehicle-to-everything (V2X) communication. We devise a simple yet effective collaboration method that achieves better bandwidth-performance tradeoffs than prior arts while minimizing changes made to single-vehicle detection models and assumptions on inter-agent synchronization. Experiments on the V2X-Sim dataset show that our collaboration method achieves 98% performance of the early collaboration while consuming the equivalent amount of bandwidth usage of late collaboration which is 0.03% of early collaboration. The code will be released at https://github.com/quan-dao/practical-collab-perception
@@
V2X Collaborative Perception, 3D Object Detection, Scene Flow.
@@ I Introduction
Object detection, which amounts to localizing objects as 3D bounding boxes, is a fundamental autonomous driving software stack module. It provides input to downstream tasks such as object tracking, motion forecast, and navigation. Due to the need for accurately localizing objects in 3D, state-of-the-art methods use LiDARs as the primary sensing modality. While significant advancements have been made, two challenges continuously plague LiDAR-based detection models: occlusion and sparsity, especially at long ranges.
As occluded objects in one frame can be visible in others due to their motion and/or the ego vehicle’s motion, and sparse regions can become dense as the ego vehicle approaches, the utilization of point cloud sequences is an appealing candidate for boosting the boundary of 3D object detection. First introduced in [2] as a means for alleviating the extreme sparsity of point clouds obtained by a 32-beam LiDAR, the concatenation of point cloud sequences using ego vehicle pose; thus the name Ego Motion Compensation (EMC), is by far the most popular multi-frame method due to its simplicity. Yet, being highly effective [3, 4, 5, 6, 7]. More recent works such as 3D-MAN [8], MoDAR [9], and MPPNet [10] are built upon this concept by dividing the input sequence into several subsequences whose point clouds are concatenated to produce preliminary detections. These preliminary detections are then fused to result in the final detections.
The major drawback of EMC-concatenated point clouds is that their formulation only accounts for the motion of the ego vehicle while ignoring the motion of dynamic objects, thus their instantiating at multiple locations, which is referred to use the shadow effect [11], during the observation horizon as can be seen in Fig.1. This effect, which is essentially a misalignment in 3D space between objects’ points and their location, results in a misalignment in feature spaces such as on the bird-eye view image, limiting the performance gain brought by an increased number of points (thanks to point cloud concatenation) to stationary and slow-moving objects [8]. In our previous work [1], we resolved this alignment explicitly by rectifying concatenated point clouds using estimated point-wise scene flow. The add-on module named Aligner, which we proposed, achieved competitive scene flow performance while substantially improving the detection accuracy, especially of large objects with a severe shadow effect.
Figure 1: A dynamic car’s appearance on the concatenation of a 10-point-cloud sequence spanning 0.5 seconds. Points are colour-coded according to their time lag with respect to the present.
Despite the power of point cloud sequences, the issue of occlusion and sparsity at long range persists as several regions remain invisible to the ego vehicle due to the geometry of the scene (e.g., a large object obstructs the ego vehicle’s field of view) or the traffic rule that limits the ego vehicle’s motion (e.g., the ego vehicle must stop at a red light). These issues can be safety critical when the ego vehicle navigates complex traffic, such as at intersections where a Waymo report indicates that a quarter of severe accidents involving their autonomous vehicles are due to occlusion [12]. Therefore, in this work, we aim to boost further object detection made by a single-vehicle operating on point cloud sequences using the Vehicle-to-everything (V2X) collaborative perception.
In the collaborative perception framework via V2X, Intelligent Road-Side Units (IRSU), which are equipped with advanced sensing systems (e.g., comparable to an autonomous vehicle) and located at elevated locations that maximize their field of view, and other vehicles can make measurements in regions that are inaccessible or partially occluded to others. Together they can come up with a complete representation of the environment. The enhanced perception capacity, thanks to V2X, comes with several new technical challenges; the most notorious among them is the performance-bandwidth tradeoff which revolves around two questions (i) which information to be broadcasted and (ii) how to fuse the exchanged information.
@@
This tradeoff establishes a design spectrum ranging from early collaboration, which exchanges raw sensory measurements - point clouds in the context of this paper, thus consuming the maximum bandwidth, to achieve the maximal performance to late collaboration that exchanges outputs (e.g., object detection as 3D bounding boxes) to achieves minimal bandwidth usage at the expense of limited performance. Naturally, researches on V2X collaboration frameworks [13, 14, 15, 16] are drawn toward the middle of this spectrum, thus the category’s name of mid-collaboration, where intermediate representations such as bird-eye view (BEV) images of agents’ surrounding environment are chosen for broadcasting in an attempt of balancing the two mutually excluding design targets. While the motivation is just, most mid-collaboration methods require making substantial changes to the architecture of single-agent perception models to accommodate the fusion module where the combination of exchanged representations takes place. More importantly, most methods make strong assumptions about the synchronization among connected agents. For example, DiscoNet [14] considers a perfectly synced setting where agents share the same clock, collect and process point clouds at the same rate and the transmission/ receiving of BEV images experience zero latency. V2VNet [13] and ViT-V2X [15] account for latency by postulating a global misalignment between exchanged BEV images and the BEV image of the ego vehicle caused by the movement of the ego vehicle between the time step it makes the requests for BEV images and the time step the exchanged BEV images are received. This means these works implicitly assume agents in the V2X network obtain point clouds synchronously.
Taking a different approach to mid-collaboration, we aim for a more practical collaboration framework that emphasizes minimal
  • Bandwidth consumption
  • Changes made to single-agent models
  • Inter-agent synchronization assumption
@@
We ground our design at minimal bandwidth consumption which can only be achieved by exchanging agents’ output which detected objects. This design choice naturally satisfies the second design target as it dismisses the need for complex mid-representation fusion modules. We achieve the final target by only assuming that connected agents share a common time reference which is practically achievable using GPS time.
The challenge posed by our relaxed inter-agent synchronization assumption is that information (i.e., detected objects) broadcasted by agents in the V2X network may never have the same timestamp as the query made by the ego vehicle. In other words, the detections made by other agents that are available on the V2X network always have an older timestamp compared to the current timestamp of the ego vehicle. This timestamp mismatch results in a misalignment between exchanged detected objects and their associated ground truths (if the detections are true positive). We resolve this misalignment by predicting objects’ velocities simultaneously with their poses using the Aligner module developed by our previous work [1]. Specifically, each detected object pools point-wise scene flow produced by the Aligner to compute its velocity, which is then used to propagate this object from one timestamp to another. As the scene flow plays a more critical role in the V2X collaborative perception, we first extend our Aligner by incorporating HD Maps into its framework and improving its supervision by distilling an Oracle model (details are in Sec.III-A2). The resulting module is dubbed Aligner++.
Aware of the limitation of the late collaboration whose details are presented in Sec.II, we derive a novel collaboration strategy that fuses detections made by connected agents at a past timestamp with the current point cloud acquired by the ego vehicle using inspiration from the recent advance in multi-frame single-vehicle object detection [9].
@@
  • Extending our previous work [1] to further boosts single-vehicle object detection accuracy and achieves more accurate scene flow prediction
  • Deriving a practical framework for V2X collaborative perception that consumes minimal bandwidth while achieving competitive performance compared to the early collaboration in the perfect synchronization setting
@@
  • Performing extensive evaluations on NuScenes [2], KITTI [17], and V2X-Sim [18] dataset to verify our method
II Related Works
A V2X cooperative perception framework comprises two main components connected autonomous vehicles (CAVs) and intelligent roadside units (IRSUs), both equipped with advanced sensing systems to perceive their environment. IRSUs are stationery installed at elevated positions to obtain minimally occluded fields of view. CAVs and IRSUs can exchange information to enhance their perception capacity (e.g., seeing beyond occlusion). In the context of this work, the sensing system of CAVs and IRSUs is restricted to a single LiDAR and the perception task focuses on 3D object detection.
@@ II Related Works
As described in the previous section, the main challenge of V2X cooperative perception is the performance-bandwidth tradeoff which establishes a design spectrum ranging from early to late collaboration. In the early collaboration approach, as depicted in Fig.2, agents exchange their raw measurements - point clouds. At every timestep, the ego vehicle concatenates its own point cloud with those obtained by other CAVs and IRSUs to form the input for its perception module. This combined point cloud offers a comprehensive view of the scene with minimal occlusion and sparsity thanks to the diverse perspectives of connected agents, thus being regarded as the upper bound of the performance of the V2X collaborative perception [18, 19]. However, due to the significant amount of bandwidth required to transmit raw point clouds (the order of 10 MB), the early collaboration strategy is not feasible for real-world deployments.
Figure 2: Architecture of the early collaboration.
@@
On the other extreme of the performance-bandwidth tradeoff, the late collaboration approach focuses on minimizing bandwidth usage by exchanging only the agents’ predictions, specifically the detection results in the form of 3D bounding boxes. As illustrated in Fig.3, each agent independently conducts object detection on its own point cloud. Subsequently, the agent merges its own predictions with those made by others to generate the final output. While this strategy is more feasible for real-world deployments thanks to its minimal bandwidth consumption, it exhibits significantly lower performance gains compared to early fusion. The limited interaction among agents in the late collaboration approach contributes to this inferior performance. In noisy environments where latency is a factor, the late collaboration strategy even underperforms single-vehicle perception [15].
Figure 3: Architecture of the late collaboration.
@@
The mid-collaboration strategy aims to find a balance between performance and bandwidth consumption by exchanging intermediate scene representations generated by the backbone of the agents’ perception model. The motivation behind this approach is that the intermediate scene representation contains more contextual information compared to the final output (i.e., 3D bounding boxes), enabling deeper interaction among agents. Moreover, this representation is more compact than raw point clouds since it has been reduced in size through a series of convolution layers in the backbone. It can also be further compressed using an autoencoder to minimize bandwidth usage. While the idea is elegant, implementing the intermediate collaboration strategy requires a range of modules, shown in Figure 4, including compressor, decompressor, and representation fusion, among others, to match the performance of early collaboration. The fusion, in particular, is quite intricate as it involves learnable collaboration graphs using techniques such as Graph Neural Networks [13, 14] or Transformers [15] to effectively fuse exchanged representations. Furthermore, dedicated modules are required to account for different practical challenges. For example, [13, 15] use the Spatial Transformer [20] to resolve the global misalignment between the ego vehicle’s representation and others’ caused by the ego vehicle’s motion between when it makes the query and when it receives exchanged information. In [16], the Fused Axial Attention [21] is used to bridge the domain gap between representations made by different detection models (e.g., PointPillar [22] and VoxelNet[23]) used by different agents in the V2X network. More importantly, most mid-collaboration methods make strong assumptions about inter-agent synchronization which is either (i) perfect synchronization where exchanged representations always share the same timestamp [14] or (ii) synchronized point cloud acquisitions, meaning all agents obtain and process point clouds at the same rate and at the same time [13, 15]. Because of these complexities, the real-world deployment of intermediate fusion remains challenging.
Figure 4: V2X Cooperative Perception via the Mid-fusion architecture.
In this work, we aim to resolve the aforementioned complexities of mid-collaboration to obtain a practical framework for V2X collaborative perception. Our design is grounded in minimal bandwidth consumption as this is nonnegotiable in real-world deployment. To achieve this design goal, we choose object detection, in the form of 3D bounding boxes, as the information to be exchanged, which is similar to the late collaboration strategy. We further relax the assumption on inter-agent synchronization to agents sharing a common time reference (e.g., GPS time) and acknowledge that agents produce detections at different rates. As a result, exchanged detections always have older timestamps compared to the timestamp of the query made by the ego vehicle, thus risking a spatial misalignment between exchanged detections and their associated ground truth (if detections are true positive). We resolve this misalignment by predicting objects’ velocity simultaneously with their locations by pooling point-wise scene flow which can be produced by integrating our Aligner++ module to any off-the-shelf BEV-based object detectors. Finally, we avoid the inferior performance of late collaboration by devising a new collaboration strategy that fuses exchanged detections with the ego vehicle’s raw point cloud for subsequent processing using its detection model rather than fusing exchanged detections with detections made by the ego vehicle. An illustration of our approach is presented in Fig.5.
The innovation of our collaboration strategy lies in recognizing the similarity between object detection using point cloud sequences and collaborative detection. In both cases, there is a need to fuse information obtained from diverse perspectives. Point cloud sequences involve capturing the motion of the ego vehicle, which results in varying viewpoints, while collaborative detection entails incorporating insights from other agents present in the environment. By drawing this parallel, we leverage the shared principle of fusing information from multiple perspectives to enhance the accuracy and robustness of both object detection approaches
Figure 5: Our V2X Cooperative Perception.
@@ III Methodology
This section first presents the extension we introduce to our previous work [1] to improve the scene flow estimation as well as further boost the performance gain when integrating our module to single-vehicle object detection models. Next, our strategy for V2X collaborative perception that enables using asynchronous exchanged information is derived.
III-A Improving Single-Vehicle Perception

### v3

@@ Abstract
Occlusion is a major challenge for LiDAR-based object detection methods as it renders regions of interest unobservable to the ego vehicle. A proposed solution to this problem comes from collaborative perception via Vehicle-to-Everything (V2X) communication, which leverages a diverse perspective thanks to the presence of connected agents (vehicles and intelligent roadside units) at multiple locations to form a complete scene representation. The major challenge of V2X collaboration is the performance-bandwidth tradeoff which presents two questions (i) which information should be exchanged over the V2X network, and (ii) how the exchanged information is fused. The current state-of-the-art resolves to the mid-collaboration approach where Birds-Eye View (BEV) images of point clouds are communicated to enable a deep interaction among connected agents while reducing bandwidth consumption. While achieving strong performance, the real-world deployment of most mid-collaboration approaches are hindered by their overly complicated architectures and unrealistic assumptions about inter-agent synchronization. In this work, we devise a simple yet effective collaboration method based on exchanging the outputs from each agent that achieves a better bandwidth-performance tradeoff while minimising the required changes to the single-vehicle detection models. Moreover, we relax the assumptions used in existing state-of-the-art approaches about inter-agent synchronization to only require a common time reference among connected agents, which can be achieved in practice using GPS time. Experiments on the V2X-Sim dataset show that our collaboration method reaches 76.72 mean average precision which is 99% the performance of the early collaboration method while consuming as much bandwidth as the late collaboration (0.01 MB on average). The code will be released in https://github.com/quan-dao/practical-collab-perception.
@@
collaborative perception, V2X, 3D object detection, deep learning, LiDAR.
@@ I Introduction
The 3D detection and localization of objects is a fundamental module of an autonomous driving software stack. It provides input to downstream tasks such as object tracking, motion prediction, and navigation. Due to the requirement of localizing objects accurately in 3D, state-of-the-art methods use LiDARs as a primary sensing modality. While significant advancements have been made, occlusion remains troublesome for LiDAR-based detection models, which makes regions of interest inaccessible to the ego vehicle. The issue of occlusion becomes particularly critical when navigating complex traffic scenarios, such as intersections, as autonomous vehicles are required to avoid collisions using a field of view that is severely reduced due to a large number of road users. The need for addressing occlusions at intersections is highlighted in a report published by Waymo [1] stating that two out of eight severe accidents involving their autonomous vehicles are due to occlusion in the intersections.
Vehicle-to-Everything (V2X) collaborative perception is a promising solution to the occlusion challenge. Its core idea is to form a complete scene representation using measurements collected from multiple perspectives by leveraging the communication among multiple connected agents presenting at different locations. Connected agents can be either connected and automated vehicles (CAVs) or Intelligent Road-Side Units (IRSUs), which are advanced sensing systems strategically placed at elevated locations to have maximal coverage of regions where complex traffic scenarios take place. This enhanced perception capacity, thanks to V2X, comes with several new technical challenges; the most notorious among them being the performance-bandwidth tradeoff which presents two questions; (i) which information should be broadcast, and (ii) how the exchanged information should be fused.
This tradeoff establishes a spectrum of solutions ranging from early to late collaboration. Raw measurements, which are point clouds in the context of this paper, are exchanged in the framework of early collaboration to maximally reduce the impact of occlusion, thus achieving the highest performance at the expense of spending a significant amount of bandwidth. On the other extreme, late collaboration exchanges high-level outputs (e.g., object detection as 3D bounding boxes) to minimize bandwidth usage, though this approach has been demonstrated to have reduced performance gains. In an attempt to balance the two mutually excluding design targets, research on V2X collaboration frameworks [2, 3, 4, 5] are drawn toward the middle of this spectrum, thus the category’s name of mid-collaboration, where intermediate representations such as BEV images of agents’ surrounding environment are chosen for broadcasting.
@@
While the motivation is just, most mid-collaboration methods require making substantial changes to the architecture of single-agent perception models to accommodate the fusion module where the combination of exchanged representations takes place. More importantly, these methods make strong assumptions about data synchronization among connected agents. For example, DiscoNet [3] considers a perfectly synced setting where agents share the same clock, collect and process point clouds at the same rate and the transmission/ receiving of BEV images experience zero latency. V2VNet [2] and ViT-V2X [4] account for latency by postulating a global misalignment between exchanged BEV images and that of the ego vehicle. The cause of this misalignment is the movement of the ego vehicle between the time step it queries the V2X network and the time when the exchanged BEV images are received. This implicitly assumes that agents in the V2X network obtain point clouds synchronously.
Another drawback of the current state-of-the-art of V2X collaborative perception is that only one point cloud per agent is used. Given that objects obscured in one frame may become visible in subsequent frames due to their movement or the motion of connected agents, and sparse regions might become dense as they draw nearer, harnessing sequences of point clouds is a compelling strategy to improve the performance of collaborative perception. In fact, the rich literature on multi-frame methods for single-vehicle object detection [6, 7, 8, 9, 10, 11] has confirmed the effectiveness of point cloud sequences as a simple concatenation of point clouds in a common frame can boost detection accuracy by approximate $30\%$ [6].
Finally, mid-collaboration methods assume connected agents share a common type of detector which is impractical for real-world deployment. This assumption is critical because the domain gap between the BEV representation of point clouds made by different detectors severely affect their fusion [5]. Prior work [5] resolves this challenge by introducing an additional module on top of those needed for the mid-collaboration to account for various differences in the BEV images made by SECOND [12] and PointPillar [13], thus further complicating the collaborative perception architecture.
Aware of the aforementioned drawbacks of previous works on V2X collaborative perception, we seek a practical collaboration framework that emphasizes:
  • Minimal bandwidth consumption
  • Minimal changes made to single-agent models
  • Minimal inter-agent synchronization assumptions
  • Support heterogeneous detectors networks
@@
We aim our design at minimal bandwidth consumption, which can only be achieved by exchanging information about the objects detected by each agent. This design choice naturally satisfies the second design target as it dismisses the need for complex mid-representation fusion modules. We achieve the third target by only assuming that connected agents share a common time reference which is practically achievable using GPS time. To reach the last target, we decide to perform the fusion of exchanged information in the input of the ego vehicle’s detector.
The challenge posed by our relaxed inter-agent synchronization assumption is that information (i.e., detected objects) broadcast by agents in the V2X network may never have the same timestamp as the query made by the ego vehicle. In other words, the detections made by other agents that are available on the V2X network always have an older timestamp compared to the current timestamp of the ego vehicle. This timestamp mismatch results in a misalignment between exchanged detected objects and their associated ground truths (if the detections are true positives), thus risking the overall performance. Our solution to this issue lies in the information that prior works have neglected - point cloud sequences. We reason that objects detected in the past can be propagated to the present if their velocities are available. The prediction of objects’ velocities pertains to motion prediction [14, 9, 15] or object tracking [16] which can negate our minimal architecture changes design target. Since we assume agents produce predictions at least at the rate they acquire point clouds, we only need short-term (e.g., 0.1 seconds if point clouds are collected at 10Hz) velocity prediction, which can be computed using scene flow, rather than long-term prediction (e.g., 3 seconds [14]) offered by motion prediction or tracking. As a result, we use the plug-in module for scene flow estimation developed by our previous work [17]. This choice effectively makes our V2X collaboration framework a multi-frame method, thus enabling each connected agent, as an individual, to enjoy a boost in detection accuracy as single-vehicle multi-frame methods do.
@@
  • Deriving a practical framework for V2X collaborative perception that outperforms Early Collaboration while consuming as much bandwidth as Late Collaboration. In addition, our method does not make any assumptions about inter-agent synchronization except the existence of a common time reference, introduces minimal changes to the architecture of single-vehicle detectors, and supports heterogeneous detector networks.
  • Demonstrate the benefit of point cloud sequences in V2X collaborative perception
@@
  • Extending our previous work [17] to further boost single-vehicle object detection accuracy and achieves more accurate scene flow prediction
@@
  • Performing extensive evaluations on NuScenes [6], KITTI [18], and V2X-Sim [19] datasets to validate the performance our method
The remainder of this paper is organized as follows. Sec.II reviews the related works on V2X collaborative perception and highlights the gap to which we aim to contribute. Next, our methodology is explained in Sec.III. Sec.IV presents our experimental results. Finally, the conclusion and future perspective are drawn in Sec.V.
II Related Work
@@
As described in the previous section, the main challenge of V2X cooperative perception is the performance-bandwidth tradeoff which establishes a solution spectrum ranging from early to late collaboration. In the early collaboration approach, as depicted in Fig.1, agents exchange their raw measurements - point clouds. At every timestep, the ego vehicle concatenates its own point cloud with those obtained by other agents to form the input for its perception model. This combined point cloud offers a comprehensive view of the scene with minimal occlusion and sparsity thanks to the diverse perspectives of connected agents, thus being regarded as the upper bound of the performance of the V2X collaborative perception [19, 20]. However, due to the significant amount of bandwidth required to transmit raw point clouds (the order of 10 MB), the early collaboration strategy is not feasible for real-world deployments.
Figure 1: Early Collaboration framework. Point clouds broadcast by connected agents are concatenated with that of the ego vehicle, forming the input to its detector.
@@
On the other extreme of the performance-bandwidth tradeoff, the late collaboration approach focuses on minimizing bandwidth usage by exchanging only the detection results from each agent in the form of 3D bounding boxes. As illustrated in Fig.2, each agent independently detects objects using its own point cloud. Subsequently, the agent merges its own predictions with those made by others to generate the final output. While this strategy is more feasible for real-world deployments thanks to its minimal bandwidth consumption, it exhibits significantly lower performance gains compared to early fusion. The limited interaction among agents in the late collaboration approach contributes to this inferior performance. In noisy environments where latency is a factor, the late collaboration strategy even underperforms single-vehicle perception [4].
Figure 2: Late Collaboration framework. Objects detected by other agents are fused with those detected by the ego vehicle to form its final output.
@@
The mid-collaboration strategy aims to find a balance between performance and bandwidth consumption by exchanging intermediate scene representations generated by the backbone of the agents’ perception model. The motivation behind this approach is that the intermediate scene representation contains more contextual information compared to the final output (i.e., 3D bounding boxes), enabling deeper interaction among agents. Moreover, this representation is more compact than raw point clouds since it has been reduced in size through a series of convolution layers in the backbone and can be further compressed using an autoencoder to minimize bandwidth usage. While the idea is elegant, implementing the intermediate collaboration strategy requires a range of modules, shown in Fig.3, including compressor, decompressor, and representation fusion, among others, to match the performance of early collaboration. The fusion, in particular, is quite intricate as it involves learnable collaboration graphs using techniques such as Graph Neural Networks [2, 3] or Transformers [4] to effectively fuse exchanged representations.
Figure 3: Mid Collaboration framework. The ego vehicle fuses the BEV image of its point cloud with those broadcast by other agents to improve its performance.
In addition, dedicated modules are required to account for different practical challenges. For example, [2, 4] use the Spatial Transformer [21] to resolve the global misalignment between the ego vehicle’s representation and others’ caused by the ego vehicle’s motion between when it makes the query and when it receives exchanged information. In [5], the Fused Axial Attention [22] is used to bridge the domain gap between representations made by different detection models (e.g., PointPillar [23] and SECOND[12]) used by different agents in the V2X network. Finally, most mid-collaboration methods make strong assumptions about inter-agent synchronization which is either (i) perfect synchronization where exchanged representations always share the same timestamp [3] or (ii) synchronized point cloud acquisitions, meaning all agents obtain and process point clouds at the same rate and at the same time [2, 4]. Because of these complexities, the real-world deployment of intermediate fusion remains challenging.
@@ III Methodology
In this work, we aim to resolve the aforementioned complexities of mid-collaboration to obtain a practical framework for V2X collaborative perception. Our design is grounded in minimal bandwidth consumption as this is essential for real-world deployment at scale. To achieve this goal we choose object detection, in the form of 3D bounding boxes, as the information to be exchanged which is similar to the late collaboration strategy. We further relax the assumption on inter-agent synchronization to agents sharing a common time reference (e.g., GPS time) and acknowledge that agents produce detections at different rates. As a result, exchanged detections always have older timestamps compared to the timestamp of the query made by the ego vehicle, thus risking a spatial misalignment between exchanged detections and their associated ground truth (if detections are true positive). We resolve this misalignment by predicting objects’ velocity simultaneously with their locations by pooling point-wise scene flow which can be produced by integrating our Aligner module [17] to any BEV-based object detectors. Finally, we avoid the inferior performance of late collaboration by devising a new collaboration strategy that fuses exchanged detections with the ego vehicle’s raw point cloud for subsequent processing by its detection model. The resulting framework is illustrated in Fig.4. We name our method late-early collaboration as connected agents broadcast their outputs, which is the signature of late collaboration, while the ego vehicle fuse received information with its point cloud, which is the signature of early collaboration.
Figure 4: Our late-early collaboration framework for V2X Cooperative Perception. In which, objects detected by each connected agent at a past time that is closest to the present are broadcast. Exchanged detected objects are propagated to the present using their velocity computed by pooling point-wise scene flow, then fused with the point cloud which the ego vehicle collected at the present to enhance its perception.
The innovation of our collaboration strategy lies in our recognition of the similarity between object detection using point cloud sequences and collaborative detection. In both cases, there is a need to fuse information obtained from diverse perspectives. Point cloud sequences involve capturing the motion of the ego vehicle, which results in varying viewpoints, while collaborative detection entails incorporating insights from other agents present in the environment. By drawing this parallel, we leverage the shared principle of fusing information from multiple perspectives to enhance the accuracy and robustness of both object detection approaches. Specifically, we utilize the latest advancement in the multi-frame object detection literature, called MoDAR [24], which interprets previously detected objects to 3D points with additional features made of object sizes, heading, confident score, and predicted class. These points are propagated to the present and merged with the point cloud obtained at the present to form the input of any off-the-shelf detectors.
@@
Since the scene flow plays a prominent role in our V2X collaboration framework by propagating past detections to the present time, this section first recalls our previous work [17] and presents the extension we make to improve the accuracy of its scene flow prediction. Next, our strategy for V2X collaborative perception that enables using asynchronous exchanged information is derived.
III-A Learning Scene Flow